a.Nellie's age is 2 years more than twice Mark's age. Last year Mark was one-third of Nellie's age. How old are Mark and Nellie?
b. Andrea had twin sisters, Emma and Ethel. In 4 years Andrea will be as old as the twins are now. The sum of the sister's age is 24. How old is each sister?

Answer by Aldorozos(172)

Equation One: N=2m+2
Equation two: 1/3 (N-1) = (m-1) multiply this equation by 3 to get rid of the denominator: (N-1) = 3m - 3
These are two equations two unknown. m is for mark and N is for Nellie
Solving the following two equations:
N=2m+2
(N-1) = 3m - 3 by subtracting the second equation from the first we get m=4 and N=10
Second problem
E for one the twin sisters
E for another twin
E-4 is for Andrea since Andrea is four years younger than twins
We add all three sister's ages: E+E+E-4 = 24
3E - 4 = 24 solving this equation we get E = 9and 1/3 = age of the twin sisters
Considering that Andrea is 4 years younger her age = 9 and 1/3 - 4 = 5 and 1/3